John is 5 more than 3 times as old as his son. If the sum of thier ages is 73 what is the age of each pearson
Answer by Jk22(389)

j=john's age, s=son's age.
j=5+3s
j+s=73 (put j from previous in this equation:)
5+3s+s=73 => 4s=68 => s=17 y, son's age
j=5+3*17=5+30+21=56 y, John's age.